A 65.0-g object connected to a spring with a force constant of 25.0 N/m oscillates with an amplitude of 7.00 cm on a frictionless, horizontal surface.

(a) Find the total energy of the system.
mJ

(b) Find the speed of the object when its position is 1.30 cm. (Let 0 cm be the position of equilibrium.)
m/s

(c) Find the kinetic energy when its position is 3.00 cm.
mJ

(d) Find the potential energy when its position is 3.00 cm.
mJ
Answer #1

a) total energy = PE of the spring at the amplitude

TE = 1/2 kx^2

TE = 0.5*25*(0.07)^2

TE = 0.06125 J = 61.25 mJ

b) speed of the object at 1.3 cm

TE = energy at 1.3

TE = PE + KE

0.06125 = 0.5mv^2 + 0.5kx^2

0.06125 = 0.5*0.065*v^2 + 0.5*25*(0.013)^2

v = 1.35 m/s

so, the velocity at 1.3 cm = 1.35 m/s

c) and d) TE = KE +PE

PE at x = 3cm = 0.5*25*(0.03)^2 = 0.01125 = 11.25 mJ

KE = TE-PE = 61.25 - 11.25 = 50mJ

c) 50 mJ

d) 11.25 mJ